# A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data

### v1

@@ Abstract
Lexical Semantic Change (LSC) offers insights into cultural and social dynamics. Yet, the validity of methods for measuring kinds of LSC has yet to be established due to the absence of historical benchmark datasets. To address this gap, we develop a novel three-stage evaluation framework that involves: 1) creating a scalable, domain-general methodology for generating synthetic datasets that simulate theory-driven LSC across time, leveraging In-Context Learning and a lexical database; 2) using these datasets to evaluate the effectiveness of various methods; and 3) assessing their suitability for specific dimensions and domains. We apply this framework to simulate changes across key dimensions of LSC (SIB: Sentiment, Intensity, and Breadth) using examples from psychology, and evaluate the sensitivity of selected methods to detect these artificially induced changes. Our findings support the utility of the synthetic data approach, validate the efficacy of tailored methods for detecting synthetic changes in SIB, and reveal that a state-of-the-art LSC model faces challenges in detecting affective dimensions of LSC. This framework provides a valuable tool for dimension- and domain-specific benchmarking and evaluation of LSC methods, with particular benefits for the social sciences.
A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data
@@ 1 Introduction
Lexical Semantic Change (LSC) provides a unique window into cultural dynamics by revealing how language evolution reflects social changes. Recently developed state-of-the-art (SOTA) computational methods have expanded our ability to classify established types of LSC, such as generalization and specialization Cassotti et al. (2024a). Efforts have also been directed towards developing methods for measuring newly proposed dimensions of LSC Baes et al. (2024); de Sá et al. (2024). Nevertheless, the field faces challenges in validating these methods. A major obstacle is the absence of historical benchmark datasets, which restricts the standardization and fair comparison of metrics. Additionally, there is a pressing need for fine-grained evaluation methods that save time and resources.
To address these challenges, the present study introduces a three-stage evaluation framework. It: 1) develops a scalable, domain-general methodology for generating high-quality synthetic sentences that leverage In-Context Learning (ICL) and a lexical database to simulate changes in kinds of LSC; 2) uses these newly constructed historical datasets to evaluate the relative effectiveness of computational approaches; and 3) identifies the more suitable method for specific dimensions and domains. This framework is applied to assess the sensitivity of various methods to detect synthetic change in major LSC dimensions–Sentiment, Intensity, and Breadth (SIB; Baes et al. 2024)–using examples drawn from psychology. Our findings confirm the validity of theory-driven changes using synthetic SIB datasets and emphasize the need to tailor methods to particular dimensions, as the SOTA LSC model was found to be ineffective at detecting affective dimensions. This framework provides an efficient and scalable solution for dimension- and domain-specific benchmarking and evaluation of LSC methods. While this innovation is generally applicable, it is particularly beneficial for the social sciences and humanities, where customized methods are essential for analyzing complex constructs.
@@ 2.1 Theoretical Background
Linguists have long debated taxonomies of LSC Bloomfield (1933); Blank (1999), defined as innovations which change the lexical meaning of a form Bloomfield (1933). A growing body of work has identified ways to detect changes in the meanings of words and quantify the extent of these changes using a variety of computational approaches (Kutuzov et al., 2018; Tahmasebi et al., 2018; Tang, 2018; Cassotti et al., 2024b; Periti and Montanelli, 2024a; Kiyama et al., 2025).
Recent years have seen the development of theoretical frameworks that propose multiple dimensions of LSC. Baes et al. (2024) introduced a three-dimensional framework that maps LSC along axes of SIB, reflecting a word’s acquisition of more positive or negative connotations (Sentiment), more or less emotionally charged or potent connotations (Intensity), and the expansion or contraction of its semantic range (Breadth). It draws on linguistic Geeraerts (2010) and psychological (Haslam, 2016) theories, and provides methodological tools to estimate SIB across time. In parallel, de Sá et al. (2024) proposed a framework that clusters LSC into three dimensions using graph structures: Orientation (shifts towards more pejorative or ameliorated senses), Relation (changes towards metaphoric or metonymic usage), and Dimension (variations between abstract/general and specific/narrow meanings). While de Sá et al. (2024) surveyed statistical methods for representing word meaning (word frequency, topic modeling, and graph structures) on dimensions, they did not demonstrate their usage.
@@ 2.2 Evaluation
The first human-annotated dataset of LSC in multiple languages (English, German, Latin, Swedish; Schlechtweg et al., 2020) represented substantial progress in indicating the presence and degree of LSC, but omitted information about kinds of change. Creating expert-annotated datasets of LSC is costly and time-intensive. Recognizing this gap, Dubossarsky et al. (2019) introduced a method to artificially induce semantic change in controlled testing environments, allowing for precise testing of how well models capture these shifts.
Recent developments in generative artificial intelligence highlight the potential of pre-trained LLMs to adapt to novel tasks at inference time through ICL Zhou et al. (2023). Few-shot ICL, a paradigm that enables LLMs to learn tasks by analogy given only a few demonstrative examples, helps to incorporate theoretical knowledge without needing to fine-tune its internal parameters Dong et al. (2024). Instead, ICL uses context from the model’s prompt to adapt the LLM to downstream tasks Radford et al. (2019); Brown et al. (2020); Liu et al. (2024). de Sá et al. (2024) demonstrated the utility of few-shot ICL, employing Chain-of-Thought and rhetorical devices, to annotate LSC dimensions, but their strategy focuses on multi-class classification of change between two sense usages. ICL offers a promising solution to bridge the absence of standardized approaches (Hengchen et al., 2021) for assessing the effectiveness of different methods to measure dimensions of LSC.
2.3 The Present Study
The present study aims to develop an evaluation framework that: (1) creates a scalable, domain-general methodology for constructing high-quality LLM-generated datasets labeling changes in LSC dimensions; (2) uses these synthetic datasets to compare the validity of proposed computational approaches; and (3) identifies the suitability of methods for each dimension and domain. We apply this framework to major LSC dimensions defined by Baes et al. (2024) (SIB; see Table 1) on a sample of words drawn from a corpus of academic psychology articles. Key questions include:
  1. 1.
    Can synthetic datasets validate methods to measure dimensions of LSC? We predict that SIB scores will be linearly associated with levels of synthetic change.
  2. 2.
    Which out of a set of LSC detection methods is most sensitive to synthetically induced changes in SIB?
Dimension Definition Examples of Rising Examples of Falling
Sentiment Relates to the degree to which a word’s meaning acquires more positive (‘elevation’, ‘amelioration’) or negative (‘degeneration’, ‘pejoration’) connotations. craftsman, once associated with manual labor, has now come to convey artistry, skill, and high-quality workmanship.
geek, from a derogatory term for odd people, to reference someone passionate about a specific field.
awful has shifted from its original meaning of "awe-inspiring" to its modern usage which indicates something very bad.
Intensity Relates to the degree to which a word’s meaning changes to acquire more (‘meiosis’) or less (‘hyperbole’) emotionally charged (i.e., strong, potent, high-arousal) connotations. cool has evolved from describing temperate to expressing strong approval or trendiness.
love, evolved from a romantic or platonic attachment to a milder expression of liking (e.g., "I love pizza.")
trauma, from referencing brain injuries to referring to less severe events (e.g., business loss).
Breadth Relates to the degree to which a word expands (‘widening’, ‘generalization’) or contracts (‘narrowing’, ‘specialization’) its semantic range. cloud, initially a meteorological term, broadened its use to reference internet-based data storage.
partner, originally referring to business co-owners, now also describes a significant other in a romantic or domestic relationship.
doctor, once referring to any scholar or teacher, now primarily refers to a medical professional.
meat, originally referred to any kind of food in Old English (‘mete’), but its meaning has narrowed to specifically denote animal flesh as food.
@@ 3.1.2 WordNet
Although other ontologies were considered, [^2] PsycNET, UMLS, DSM-5, ConceptNet the English WordNet lexical database 3.0 Miller (1992) was chosen for its linguistic coverage and lexical structure. It organizes words into synsets (synonyms with distinct meanings), linking them by semantic relationships (e.g., hypernyms, hyponyms).
@@ 3.1.3 Targets
While the evaluation framework is general in its applicability, six terms from psychology—abuse, anxiety, depression, mental health, mental illness, and trauma—are analyzed for semantic change, selected for their empirical and theoretical relevance to shifting word meanings. Trauma, mental health, and mental illness have seen falls in their average valence and semantic expansions (trauma: Baes et al., 2023; Haslam et al., 2021; mental health, mental illness: Baes et al., 2024). There have been changes in the intensity of their meanings, with rises for mental health and mental illness (Baes et al., 2024), as well as anxiety and depression (Xiao et al., 2023), and a fall for trauma (Baes et al., 2023). Qualitatively, abuse has expanded horizontally to include passive neglect and emotional abuse, beyond its physical scope Haslam (2016). Targets were sufficiently prevalent (sentence counts: 46,272; 104,486; 115,430; 44,130; 5,808; 23,187). Appendix A shows annual counts.
@@ 3.2 Evaluation Framework
The general pipeline for the evaluation framework is shown in Figure 1. Synthetic datasets are constructed to benchmark changes in LSC dimensions using few-shot ICL and a lexical database. GPT-4o Achiam et al. (2023) [^3] ChatGPT API documentation: https://platform.openai.com/docs/guides/text-generation is prompted with expert-crafted examples to increase and decrease corpus sentences in affective dimensions across 5-year intervals. This ensures that synthetic sentences are theory-driven, domain-specific and contain temporal features. GPT is used due to its adeptness at few-shot learning, task adaptation with minimal examples (Achiam et al., 2023; Merx et al., 2024) and lack of disciplinary bias Ziems et al. (2024). Appendix B details the synthetic datasets, [^4] Link to Synthetic datasets: [MASKED LINK]. validated using tools that measure SIB Baes et al. (2024).
@@
To assess the relative effectiveness of different methods, sentences are sampled from the natural and synthetic corpora using two sampling strategies. Bootstrap sampling draws 50 sentences with replacement from both corpora 100 times (i.e., iterations). Five-year random sampling selects up to 50 sentences from fixed intervals 10 times, ensuring each time period is equally represented. Each iteration forces unique sentence selection while permitting sentence repetition across different rounds to reflect natural language. Control conditions shuffle these sentences to balance authentic:synthetic sentences in each sample, verifying the synthetic effect in the genuine condition and its absence in the control. Following computational linguistics precedents Dubossarsky et al. (2017, 2019), this approach validates the impact of synthetic interventions by providing a baseline for comparison. Notably, for each strategy, synthetic sentences are injected into natural samples at increasing injection levels (20%, 40%, 60%, 80%, and 100%), as illustrated in Appendix B. Bins are injection levels (for bootstrap) and epoch (for 5-year intervals). This simulates controlled semantic saturation scenarios to assess the sensitivity of methods to semantic variation (increased/decreased SIB) (stage 2) and select the method that detects greater magnitude of change from 0% to 100% injection (stage 3).
3.3 Sentiment and Intensity
3.3.1 Synthetic Sentiment and Intensity
To generate the synthetic Sentiment and Intensity datasets, we employ few-shot ICL with GPT-4o to vary these dimensions. First, neutral sentences from the corpus (detailed in Section 3.1.1) are sampled as outlined in Appendix C. Second, a psychology scholar crafts five Chen et al. (2023) diverse examples of sentence variations for each target following the task detailed below, which includes construct definitions to generate theory-driven change. For ‘scholar-in-the-loop’ few-shot demonstrations, see Appendix D (Sentiment) and Appendix E (Intensity). Third, the prompt is refined during pilot tests (10 inputs). Fourth, for each of the neutral sentences (Sentiment: 36,151; Intensity: 39,896), we make one inference call to GPT-4o through the OpenAI API to generate variations of Sentiment (positive/negative) or Intensity (high/low). Fifth, output sentences are manually adjusted (Sentiment: 0.25%; Intensity: 0.01%) due to GPT-4o’s failure to retain targets. See Appendix C for the counts of input/output sentences and final prompts (dataset costs for Sentiment: 115 \$US; Intensity: 136 \$US).
3.3.2 Quantifying Sentiment and Intensity
where $w_{i,j,k}$ denotes the frequency of each collocate $i$ in iteration $k$ within bin $t_{j}$, and $x_{i,j,k}$ denotes its valence or arousal rating at bin $t_{j}$ within iteration $k$. Here, $n_{j,k}$ is the number of collocates in iteration $k$ within bin $t_{j}$. Scores are weighted by the collocate’s frequencies within each iteration and normalized by the total occurrences in that iteration. Scores are averaged across all iterations within each bin, conditioned on whether the Sentiment is positive/negative, or the Intensity is high/low. These indices provide a mean valence or arousal score per iteration in each bin $t_{j}$, with higher scores indicating a more positive valence or higher arousal. Scores (1-9) are normalized to range from 0 (extremely unhappy/low arousal) to 1 (extremely happy/high arousal).
[…]
3.4 Breadth
3.4.1 Synthetic Breadth
Unlike Sentiment and Intensity, current Breadth measures have no score that assigns a mid-point with which to obtain neutral sentences to vary. Therefore, to simulate semantic breadth, we adapt Dubossarsky et al.’s (2019) replacement strategy, using WordNet 3.0 to expand a target word’s usage by incorporating contexts from donor terms, broadening its semantic range without altering its core meaning. Relevant synsets are identified and filtered for psychological relevance using keyword matching [^8] Psychology key terms: ”abnormality”, ”abnormally”, ”emotional”, ”feeling”, ”feelings”, ”harm”, ”hurt”,”mental”, ”mind”, ”psychological”, ”psychology”, ”psychiatry”, ”syn- drome”, ”therapy”, ”treatment”. and semantic similarity thresholds. Donor terms (co-hyponyms with the target) are filtered using Lin similarity (0.5) [^9] Information content values from the psychology corpus and cosine similarity (0.7) with embeddings from BioBERT Lee et al. (2020), a pre-trained language model for biomedical text mining, to capture context-dependent meanings of synset glosses in 768-dimensional vectors. See Appendix F for the list. The sibling replacement process identifies and replaces sibling terms with the target, shown below. To sample representatively from the sibling list, a round-robin strategy is used, sampling up to 1,500 unique sentences per epoch per injection level to create the final synthetic breadth dataset (0 \$US).
Dataset Creation for Synthetic Breadth Replacement Strategy: Randomly sample sentences containing co-hyponyms of the target term from the validated list and replace the co-hyponym with the target to be used as a synthetic sentence. ————————————————————— [One example for mental_health below.] Donor Context: The ‘Angry and Impulsive Child’ and ’Abandoned and Abused Child’ modes uniquely predicted dissociation scores. Synthetic Context: The ‘Angry and Impulsive Child’ and ’Abandoned and Abused Child’ modes uniquely predicted mental_health scores.
3.4.2 Quantifying Breadth
To estimate the semantic broadening (expansion) or narrowing (contraction) of a word’s meaning, we calculate the average cosine distance between sentence-level embeddings of a target term, as in Baes et al. (2024). The SentenceTransformer model ‘all-mpnet-base-v2’ [^10] Microsoft pretrained network (109M model params) https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is used to generate these embeddings. The Breadth score, $B$, is derived by averaging the cosine distances, $\delta$, across all unique pairs of sentence embeddings within each iteration, and then averaging these scores across all iterations within each bin, as shown in Equation 2:
Here, $\delta(s_{i,k}^{t_{j}},s_{j,k}^{t_{j}})$ calculates the cosine distance between two sentence embeddings in the same iteration $k$ in bin $t_{j}$. $N_{k}$ is the number of sentence embeddings in iteration $k$; $I_{j}$ is the number of iterations in bin $t_{j}$. Higher scores indicate greater variation in the target’s semantic range. Scores range from 0 (no variation) to 1 (max variation).
We compare the sentence transformer "all-mpnet-base-v2" (MPNet) with Cassotti et al.’s (2023) SOTA word transformer "XL-LEXEME" [^11] XL-LEXEME (~550M model params) https://huggingface.co/pierluigic/xl-lexeme (XLL). While MPNet generates sentence embeddings through pooling tokens, which dilutes word-specific information, XLL uses a bi-encoder architecture that focuses on word-specific attention, [^12] Only the first occurrence of the target is attended to. using polysemy as a proxy for meaning divergence during training (WIC; Pilehvar and Camacho-Collados, 2019).
3.5 General Lexical Semantic Change
To quantify general LSC, we use the SOTA LSC score Cassotti et al. (2023), which calculates the Average Pairwise Cosine Distances (Giulianelli et al., 2020) between sentence embeddings from two time periods. We extend it to compare embeddings from different bins within the same iteration, as shown in equation 3:
4 Results
Synthetic Change Effects:
The hypothesis that scores from Baes et al.’s (2024) SIB tools will be linearly associated with levels of synthetic change is supported, as evidenced by rising or falling trends in SIB scores across all targets and conditions (Figure 2). Mixed linear models demonstrate increases or decreases on SIB scores for every 1-unit increase in synthetic injection level (detailed in Table 2 and Appendix G). SIB scores for the five-year sampling experiments depict similar trends in response to varying injection levels (Appendix H).
Score Valence Arousal Breadth
$\beta^{+}$003* .002* <.0001*
$\beta^{-}$.001* -.002* N/A
Table 2: Coefficients of Mixed Linear Models Predicting SIB Scores from Injection Levels (Var. Target Intercept)
[…]
Control Experiments:
As illustrated in Figure 2, controlling for synthetic injection level by re-analyzing data with shuffled sentences for uniform distribution reveals flat SIB score trends in bootstrapped settings. Appendix H shows that, even with temporal shuffling within time bins, SIB scores in five-year samples tend to converge to a midpoint between natural and synthetic data.
Comparative Method Evaluation:
Comparisons of the relative validity of alternative change detection methods yielded mixed results. To determine which method is more sensitive to synthetically induced changes in SIB, we compare their performance on a synthetic change detection task using an evaluation metric specified below.131313 For XLL’s LSC Score, $\Delta$ is normalized against the intrinsic within-bin variability in both bins of interest: $\Delta\ =\frac{\text{APD}(X_{100}\text{-between-}X_{0})}{\text{max}\left[\text{APD}(X_{0}\text{-within-}X_{0}),\text{APD}(X_{100}\text{-within-}X_{100})\right]}$
Percent Relative Change Index Synthetic Change Detection Task: Detect the magnitude of change in the target word’s context when sampling sentences from a natural corpus (0%) and an entirely synthetic corpus (100%). Percent relative change $\Delta$ is defined as: $\Delta\%=\frac{X_{100}-X_{0}}{X_{0}}\times 100$ where $X_{0}$ represents the measure’s score at 0% synthetic injection (natural data) and $X_{100}$ at 100% synthetic injection (fully artificial data).
For Sentiment, Valence index and ABSA’s Sentiment score are sensitive to detecting variations in synthetic Sentiment, although the ABSA score outperforms the Valence index 10/12 times. For Intensity, the Arousal index shows sensitivity to detecting variations in synthetic Intensity. For Breadth, XLL outperforms MPNet (4/6 times) on detecting rises in synthetic Breadth using the Breadth score.
Critically, XLL-LSC score is completely insensitive to detecting changes in either Sentiment or Intensity. XLL-LSC can only indicate change via positive change values, while negative values indicate that the within-bin variance is greater than the change scores between bins. See Appendix I for between- and within-bin LSC scores across all synthetic injection levels. Thus, the negative scores observed in Sentiment and Intensity (except for Mental Illness) establish that XLL was unable to detect any change signal in these words. XLL-LSC detects changes in synthetic Breadth for 2/6 terms.
5 Discussion
The present study introduced a three-stage general domain evaluation framework that: 1) creates synthetic datasets featuring ‘scholar-in-the-loop’ LLM-generated sentences to simulate various kinds of LSC; 2) leverages these datasets to assess the sensitivity of computational approaches to synthetic changes; and 3) evaluates the suitability of these methods for specific dimensions and domains. This framework is applied to generate synthetic datasets that induce changes across the three dimensions of a recent multidimensional LSC framework (SIB; Baes et al., 2024), using examples from psychology, to test and compare the suitability of different methods in detecting these synthetic changes.
Our findings support the hypothesis that recently proposed methods (Valence index, Arousal index, Breadth score; Baes et al., 2024) detect synthetic changes on the SIB dimensions. Control analyses, which adhered to computational linguistics standards Dubossarsky et al. (2017, 2019), confirmed the absence of these effects in shuffled controls. The implications of these findings are two-fold. The ability of SIB methods to detect changes when introducing silver-label synthetic data validates their sensitivity and reliability in detecting and measuring variations in SIB, even in controlled, artificial environments. This validates the LLM-generated sentences in our ICL evaluation suites.
We demonstrated how a synthetic change detection task can assess the sensitivity of various computational approaches, guiding the selection of the most suitable model for specific dimensions and domains. Baes et al.’s (2024) tools, which validated the synthetic SIB datasets, were supported by alternative methods that consistently detected synthetic changes in SIB across all conditions and targets, providing further validation. The Valence index and Sentiment score (ABSA) identified variations in synthetic Sentiment, while Breadth scores (XLL and MPNet) detected increases in synthetic Breadth. Results suggest that these NLP-based methods are more sensitive in detecting synthetic changes than Warriner-based methods, which rely on Valence and Arousal ratings. Future empirical studies on Sentiment and Breadth may consider adopting these NLP models, either as replacements for, or in addition to, existing methods.
Notably, when computing the general LSC score using the SOTA LSC model XLL Cassotti et al. (2023), it was not sensitive to detecting Sentiment and Intensity. Although XLL shows some sensitivity to identifying synthetic increases in Breadth, it registers a more substantial change when the Breadth score is adjusted according to the method introduced by Baes et al. (2024). It uses the within-bin average cosine distance of target containing sentences as a proxy for the expansion (broadening) or contraction (narrowing) of a word’s contextual usage. The inability of XLL to detect the affective dimensions of LSC highlights the necessity of evaluating SOTA models before deploying them in new domains. Future research should investigate whether this weakness in detecting affective dimensions is specific to XLL or extends to other contextualized models in more corpora. This inquiry is particularly salient given recent advances in analyzing fine-grained, continuous semantic shifts through “diachronic word similarity matrices using fast and lightweight word embeddings over arbitrary time periods" Kiyama et al. (2025).
6 Conclusion
The current study introduced a novel general domain evaluation framework. Its three-stage pipeline involves: 1) developing a scalable methodology for generating LLM-based synthetic datasets with silver labels that simulate changes in kinds of LSC; 2) using these datasets to evaluate the relative sensitivity of computational approaches in a synthetic change detection task; and 3) identifying the most suitable method for detecting synthetically induced changes across specific dimensions and domains. We applied this framework to a set of psychological terms. Findings not only supported the validity of proposed computational methods for measuring changes in SIB, but also established a controlled experimental standard for rigorously evaluating existing LSC detection methods and exploring alternative computational approaches. This work is crucial for addressing the substantial gap created by the lack of historical benchmark datasets, which has previously hindered the standardization of metrics and fair comparison of methods. While this innovation benefits all disciplines (e.g., biomedicine, law, theology), it is particularly valuable in the social sciences and humanities, where unique methods are often required to measure complex constructs.
Limitations
Limitations inform future directions. Evaluating the quality of LLM prompt and demonstration examples in the few-shot ICL paradigm is challenging. As LLM evaluation standards are developed Chang et al. (2024); Ziems et al. (2024), future research might explore automated strategies such as updating prompts based on examples (DSPy) [^14] https://dspy.ai or comparing LLM output from different prompts using a free, unified interface. [^15] https://github.com/marketplace/models/azure-openai/gpt-4o/playground LLM choice in the evaluation pipeline could be expanded to include open-source models (e.g., FlanT5-XL, Mistral-7B, Mixtral-8x7B), enhancing its accessibility.
Furthermore, our study benefited from using GPT-4o, which is trained on US English and is therefore well-suited for analyzing texts within the Western-centric domain of psychology. However, the cultural and linguistic biases of LLMs may pose challenges for adapting our evaluation pipeline to other languages Havaldar et al. (2023), although few-shot ICL has proven effective in low-resource languages Cahyawijaya et al. (2024). Despite the tendency of LLM training data to skew towards the recent past, it successfully generated high-quality sentences that spanned a 1970 to 2019 time period. Future research should focus on refining these models to support broader application across various cultural contexts, languages, and historical periods.
Furthermore, the present study does not specify which sense of the term is semantically expanded. Attempting to integrate senses into the synthetic data generation pipeline may provide richer insights. While the specialized psychology corpus and target words exhibit limited senses, general domain corpora introduce ambiguous contexts (e.g., economic sense of “depression"). Notably, current methods for word sense disambiguation may not integrate with distributional approaches as historical linguists do not treat LSC as a set of senses.
Ethical Considerations
Acknowledgments
We acknowledge the support of the University of Melbourne’s general-purpose High Performance Computing system, Spartan Lafayette et al. (2016), which provided ample computational power to facilitate the efficient encoding of embeddings using transformer models and corpus preprocessing.
This research was supported by an Australian Government Research Training Program Scholarship and Australian Research Council Discovery Project DP210103984.
Appendix A Corpus Counts of Target Terms
Appendix B Synthetic Dimension Datasets: Details
@@
Breadth Abuse NA 5,221 (27) NA 0
Anxiety NA 13,635 (26) NA 0
Depression NA 14,463 (27) NA 0
Mental Health NA 14,638 (26) NA 0
Mental Illness NA 14,639 (26) NA 0
Trauma NA 14,650 (26) NA 0

### v2

# LSC-Eval: A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data

@@ Abstract
Lexical Semantic Change (LSC) provides insight into cultural and social dynamics. Yet, the validity of methods for measuring different kinds of LSC remains unestablished due to the absence of historical benchmark datasets. To address this gap, we propose LSC-Eval, a novel three-stage general-purpose evaluation framework to: (1) develop a scalable methodology for generating synthetic datasets that simulate theory-driven LSC using In-Context Learning and a lexical database; (2) use these datasets to evaluate the sensitivity of computational methods to synthetic change; and (3) assess their suitability for detecting change in specific dimensions and domains. We apply LSC-Eval to simulate changes along the Sentiment, Intensity, and Breadth (SIB) dimensions, as defined in the SIBling framework, using examples from psychology. We then evaluate the ability of selected methods to detect these controlled interventions. Our findings validate the use of synthetic benchmarks, demonstrate that tailored methods effectively detect changes along SIB dimensions, and reveal that a state-of-the-art LSC model faces challenges in detecting affective dimensions of LSC. LSC-Eval offers a valuable tool for dimension- and domain-specific benchmarking of LSC methods, with particular relevance to the social sciences.
[…]
LSC-Eval: A General Framework to Evaluate Methods for Assessing Dimensions of Lexical Semantic Change Using LLM-Generated Synthetic Data
@@ 1 Introduction
Lexical Semantic Change (LSC) provides a unique window into cultural dynamics by revealing how language evolution reflects social changes (McGillivray, 2020; Schlechtweg et al., 2020). Recently developed state-of-the-art (SOTA) computational methods have expanded our ability to classify established types of LSC, such as generalization and specialization Cassotti et al. (2024a). Efforts have also been directed towards developing methods for measuring newly proposed dimensions of LSC Baes et al. (2024); de Sá et al. (2024). Nevertheless, the field faces challenges in validating these methods. A major obstacle is the absence of historical benchmark datasets, which restricts the standardization and fair comparison of metrics. Additionally, there is a pressing need for fine-grained evaluation methods that save time and resources.
To address these challenges, the present study proposes LSC-Eval, a novel three-stage evaluation framework for assessing LSC across dimensions and domains. It: (1) develops a scalable, general-purpose methodology for generating high-quality synthetic sentences using ‘scholar-in-the-loop’ In-Context Learning (ICL) and a lexical resource to simulate targeted kinds of LSC; (2) uses these datasets to evaluate the validity of alternative LSC detection methods; and (3) identifies the most suitable methods for capturing change in specific dimensions and domains. We apply LSC-Eval to test the sensitivity of various methods to synthetically induced changes along three major LSC dimensions — Sentiment, Intensity, and Breadth (SIB) — as proposed in the SIBling framework (Baes et al. 2024; see Table 1), drawing from psychology examples. Illustrative cases include awesome, which has risen in sentiment (from serious to enthusiastic), declined in intensity (from awe-inspiring to casually pleasing), and broadened in scope (from solemn reverence to general positivity); trauma, which has shifted from physical injury to a broader and milder range of psychological harms; and sick, which has risen in sentiment (in slang), broadened in usage (from a medical term to a general expression of praise), and declined in intensity (from a descriptor of serious illness to one used in casual or enthusiastic contexts).
Dimension Definition Examples of Rising Examples of Falling
Sentiment Relates to the degree to which a word’s meaning acquires more positive (‘elevation’, ‘amelioration’) or negative (‘degeneration’, ‘pejoration’) connotations. craftsman, once associated with manual labor, has come to convey artistry, skill, and high-quality workmanship.
geek, originally a derogatory term for odd people, now refers to someone passionate about a field.
awful, once meaning "awe-inspiring," now indicates something very bad.
Intensity Relates to the degree to which a word’s meaning changes to acquire more (‘meiosis’) or less (‘hyperbole’) emotionally charged (i.e., strong, potent, high-arousal) connotations. cool has evolved from describing temperature to expressing strong approval or trendiness.
love has evolved from denoting romantic or platonic attachment to also expressing milder forms of liking (e.g., “I love coffee.”).
trauma has shifted from referring to brain injuries to encompassing milder events (e.g., business loss).
Breadth Relates to the degree to which a word expands (‘widening’, ‘generalization’) or contracts (‘narrowing’, ‘specialization’) its semantic range. cloud, initially a meteorological term, broadened in usage to refer to internet-based data storage.
partner, which once referred narrowly to a business associate, has broadened to denote a romantic or domestic companion.
doctor, once used for any scholar or teacher, now primarily denotes a medical professional.
meat, which originally referred to any kind of food in Old English (mete), has since narrowed to denote animal flesh as food.
The present study addresses two key research questions: (RQ1) Can synthetic datasets be used to validate methods for measuring LSC dimensions? We hypothesize that SIB scores will be associated with levels of induced change; (RQ2) Which of several LSC detection methods is most sensitive to synthetically induced changes in SIB? Our findings confirm the validity of theory-driven changes — that is, changes systematically injected into corpora based on the SIBling framework Baes et al. (2024), which targets the dimensions of Sentiment, Intensity, and Breadth. These changes were operationalized using ICL and a lexical resource to simulate diachronic semantic shifts. Results highlight the need to tailor detection methods to specific dimensions, as the state-of-the-art LSC model failed to detect affective change. The proposed framework, LSC-Eval, offers an efficient and scalable approach for dimension- and domain-specific benchmarking of LSC methods. While broadly applicable, LSC-Eval is especially valuable in the social sciences and humanities, where capturing nuanced conceptual change is essential.
@@ 2.1 Theoretical Background
Linguists have long debated taxonomies of LSC Bréal (1900); Blank (1999), defined as innovations which change word meanings Bloomfield (1933). A growing body of work has identified ways to detect changes in the meanings of words and quantify the extent of these changes using computational approaches (Kutuzov et al., 2018; Tahmasebi et al., 2018; Tang, 2018; Tahmasebi and Dubossarsky, 2023; Cassotti et al., 2024b; Periti and Montanelli, 2024a; Kiyama et al., 2025).
Recent years have seen the development of theoretical frameworks that propose multiple dimensions of LSC. Baes et al. (2024) introduced SIBling, a three-dimensional framework that maps LSC along axes of SIB, reflecting a word’s acquisition of more positive or negative connotations (Sentiment), more or less emotionally charged or potent connotations (Intensity), and the expansion or contraction of its semantic range (Breadth). It draws on linguistic Geeraerts (2010) and psychological (Haslam, 2016) theories, and provides methodological tools to estimate SIB across time. In parallel, de Sá et al. (2024) proposed a framework that clusters LSC into three dimensions using graph structures: Orientation (shifts towards more pejorative or ameliorated senses), Relation (changes towards metaphoric or metonymic usage), and Dimension (variations between abstract/general and specific/narrow meanings). While de Sá et al. (2024) surveyed statistical methods for representing word meaning (word frequency, topic modeling, and graph structures) on dimensions, they did not demonstrate their usage.
@@ 2.2 Evaluation
The first human-annotated dataset of LSC in multiple languages (English, German, Latin, Swedish; Schlechtweg et al., 2020) marked substantial progress in identifying the presence and degree of LSC but omitted information about kinds of change. However, expert-annotated datasets are costly and time-intensive to create. To address this gap, Dubossarsky et al. (2019) introduced a method to artificially induce semantic change in controlled experimental settings, enabling precise testing of how well models capture these shifts.
Recent developments in generative artificial intelligence highlight the potential of pre-trained LLMs to adapt to novel tasks at inference time through ICL Zhou et al. (2023). Few-shot ICL, a paradigm that enables LLMs to learn tasks by analogy given only a few demonstrative examples, helps to incorporate theoretical knowledge without needing to fine-tune its internal parameters Dong et al. (2024). Instead, ICL uses context from the model’s prompt to adapt the LLM to downstream tasks Radford et al. (2019); Brown et al. (2020); Liu et al. (2024). de Sá et al. (2024) demonstrated the utility of few-shot ICL, employing Chain-of-Thought and rhetorical devices, to annotate LSC dimensions, but their strategy focused on multi-class classification of change between two sense usages. ICL offers a promising solution to bridge the absence of standardized approaches (Hengchen et al., 2021) for assessing the effectiveness of different methods to measure dimensions of LSC.
@@ 3.1.2 WordNet
Although other ontologies were considered, [^2] PsycNET, UMLS, DSM-5, ConceptNet the English WordNet lexical database 3.0 Miller (1992) was chosen for its linguistic coverage and lexical structure. It organizes words into synsets, linking them by semantic relationships (e.g., hypernyms, hyponyms).
@@ 3.1.3 Targets
While the evaluation framework is general in its applicability, six terms from psychology — abuse, anxiety, depression, mental health, mental illness, and trauma — were analyzed for semantic change, selected for their empirical and theoretical relevance to shifting word meanings. Trauma, mental health, and mental illness have seen falls in their average valence alongside their semantic expansions (trauma: Baes et al., 2023; Haslam et al., 2021; mental health, mental illness: Baes et al., 2024). There have been changes in their semantic intensity, with rises for mental health and mental illness (Baes et al., 2024), as well as anxiety and depression (Xiao et al., 2023), and a fall for trauma (Baes et al., 2023). Qualitatively, abuse has expanded horizontally to include passive neglect and emotional abuse, beyond its physical scope Haslam (2016). Targets were sufficiently prevalent (sentence counts: 46,272; 104,486; 115,430; 44,130; 5,808; 23,187). Appendix A shows annual counts.
@@ 3.2 Evaluation Framework
The three general stages of the evaluation framework are illustrated in Figure 1.
@@
In Stage 1, synthetic datasets (i.e., datasets exhibiting simulated variations in linear semantic change along a specific dimension) were created to benchmark changes in LSC dimensions using few-shot ICL and a lexical database. GPT-4o Achiam et al. (2023) [^3] ChatGPT API documentation: https://platform.openai.com/docs/guides/text-generation was prompted with expert-crafted examples to increase and decrease corpus sentences in affective dimensions across five-year intervals. This ensures that synthetic sentences are theory-driven, domain-specific and contain temporal features. GPT was used due to its adeptness at few-shot learning, task adaptation with minimal examples (Achiam et al., 2023; Merx et al., 2024) and lack of disciplinary bias Ziems et al. (2024). Table 2 provides descriptive statistics for the synthetic datasets, [^4] Link to Synthetic datasets: https://github.com/naomibaes/Synthetic-LSC_pipeline which are validated using tools that have been shown to measure SIB in historical (domain-general and -specific) corpora Baes et al. (2024).
@@
Breadth abuse 5,221 (27) 0
anxiety 13,635 (26) 0
depression 14,463 (27) 0
mental health 14,638 (26) 0
mental illness 14,639 (26) 0
trauma 14,650 (26) 0